OpenStudy (anonymous):

If y = 4x -2, find y when x = 4. A. 4 B. 6 C. 14 D. 100 + 2w

OpenStudy (anonymous):

plug in 4 y = 4(4) - 2 = 16 - 2 = 14 C

OpenStudy (anonymous):

c